OpenStudy (anonymous):

Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = 3x2 - 1

OpenStudy (anonymous):

A function \(f(x)\) is even if \(f(-x)=f(x)\), and odd if \(f(-x)=-f(x)\). (Neither if neither condition is met. So to show the given function is even or odd, you plug in \(-x\). You get \[f(-x)=3(-x)^2-1=3x^2-1=f(x)\] hence \(f(x)\) is even.

OpenStudy (anonymous):

Thank you, one more question relating to this. Is this related to function inverses?